OpenStudy (anonymous):

simplify 3x X 2x squared simplify 3x X 2x squared @Mathematics

OpenStudy (anonymous):

whats big x

OpenStudy (anonymous):

multiply

OpenStudy (asnaseer):

assuming I have read the question correctly, it is:\[(3x\times 2x)^2=(6x^2)^2=36x^4\]

OpenStudy (asnaseer):

or is it:\[3x\times 2x^2=6x^3\]

OpenStudy (anonymous):

the second one is correct

OpenStudy (asnaseer):

you can use brackets to help explain the problem better

OpenStudy (asnaseer):

e.g. what is 3x X 2(x squared) but the ideal way to enter questions of this type is to use the equation editor.

OpenStudy (anonymous):

the question is simplify 3x times 2x squared